OpenStudy (anonymous):

can i get help solving f(x)=10-x, find inverse f-1(x)_______

OpenStudy (anonymous):

\[x=10-y\] solve for y you get \[y=10-x\] so \[f^{-1}(x)=10-x\] it is it's own inverse

OpenStudy (anonymous):

you can check. for example \[f(3)=10-3=7\] and \[f(7)=10-7=3\] so it is its own inverse

OpenStudy (anonymous):

so what is the answer x+10?

OpenStudy (anonymous):

can u show me the step by step process?